OpenStudy (anonymous):

The recursive formula for a sequence is Un=Un-1.+12, where Un is the nth term of the sequence and U0=7. Which explicit formula can be used to determine the nth term of the sequence? A. 7n+19 B. 7n+12 C. 7+19n D. 7+12n

OpenStudy (loser66):

\[U_1=U_0+12\\U_2=U_1+12=~~~U_0+12+12~=U_0+2*12\\U_3=U_2+12=U_0+2*12+12=U_0+3*12\\U_4=U_3+12= U_0+3*12+12=U_0+4*12\].................................................. therefore, \(U_n= U_0+n*12=7+12n\)